Which one of the following is the highest temperature?


The boiling point of water


220°F


373 K


100°C


All of the above are identical temperatures
1
Expert's answer
2019-04-03T05:02:19-0400

The highest temperature is 220°F.


The boiling point of water is 100°C.

 

tC = TK – 273 = 373K – 273 = 100⁰C.

 

tC = (TF – 32) × 5/9 = (220°F–32) × 5/9 = 104.44⁰C.

 

So 220°F is the highest temperature.